OpenStudy (anonymous):

If g(x) = 3(x − 2), find the value of x if g(x) = 6. 12 2.6 8 4

OpenStudy (anonymous):

3(x − 2) = 6 (x − 2)=2 x=4

OpenStudy (anonymous):

Same concept as before! Plug in 6 for x. Focus on getting the basic properties down.

OpenStudy (anonymous):

@staceyyyy1 what module are you on? I'm also in FLVS

OpenStudy (anonymous):

@AJ17 : they have gives the g(x) value!

OpenStudy (anonymous):

i got 12?? and module 3. 3.06

OpenStudy (anonymous):

@DHASHNI Sorry, read the question wrong. Set 3(x-2) equal to 6.

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

yepp it was 4.